# Scalable and Robust Multiband Modeling of AGN Light Curves in Rubin-LSST

### v1

@@ 3.1.1 Light Curve Simulation
We adopt the DRW model to simulate AGN light curves in individual LSST bands. We start by simulating DRW light curves in each LSST band with a 30-second cadence; this 30-second cadence is chosen to match the 30-second exposure time of each LSST single visit (Ivezić et al., 2019). The input DRW parameters across different bands are correlated according to the wavelength-dependent power-law relations presented in MacLeod et al. (2010). More specifically,
@@
where $\lambda$ is the effective wavelength of the LSST filters (Ivezić et al., 2019). We then simulate continuum lags by offsetting light curves in redder band in time with respect to light curves in bluer bands. Finally, we down-sample the simulated 30-second cadence light curve to the observing cadence of the WFD survey.
@@
where $\sigma_{\mathrm{sys}}$ is the systematic photometric error and $\sigma_{\mathrm{rand}}$ is the random photometric error. LSST aims to achieve a $\sigma_{\mathrm{sys}}<0.005$ mag (Ivezić et al., 2019); we set $\sigma_{\mathrm{sys}}$ to $0.005$ mag in our simulation. The random photometric error ($\sigma_{\mathrm{rand}}$) is expected to scale as
@@
The observing cadence and the expected $m_{5}$ at any location on the sky are predicted by LSST operation simulations (OpSims) using MAF (Marshall et al., 2017). In this work, we adopt the operation simulation version baseline_v4.0_10yrs. All our simulated AGNs also assume a fiducial $g$-band mean magnitude of 21 ($\sim 1\%$ photometric accuracy) and mean uugg and ggrr colors of 0.1 mag (Richards et al., 2001; Temple et al., 2021). An example light curve from the first set is shown in Figure 1.
@@
The first set aims to evaluate EzTaoX’s ability to recover interband lags (see Section 3.1.2), and the light curves consist of simulated LSST observations in $g$ and $r$ bands. The ggrr lag for each multiband light curve is uniformly sampled between 0.01 day and 10 days, while the input $g$-band $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ are fixed at fiducial values of 100 days and 0.112 mag, respectively. These choices of $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ are representative of those measured from typical quasar (i.e., high luminosity AGN) UV/optical light curves (e.g., Vanden Berk et al., 2004; MacLeod et al., 2010).
@@ 3.1.2 Recovery of Interband Lag
We run EzTaoX on the first set of simulated light curves to extract the lag between the $g$- and $r$-band ($\tau_{\rm gr}$), after subtracting the mean from each single-band light curve. The EzTaoX light curve model includes seven parameters: $\tau_{\rm DRW}$, $\sigma_{\mathrm{DRW}}$, $S_{\rm g}$, $S_{\rm r}$, $\mu_{\rm g}$, $\mu_{\rm r}$, and $\tau_{\rm gr}$. Here, $S_{\rm g}$ and $S_{\rm r}$ are the GP kernel scaling factors for $g$- and $r$-band observations (see Equation 5), while $\mu_{\rm g}$ and $\mu_{\rm r}$ are the mean magnitudes of the $g$- and $r$-band light curves, respectively. As part of the EzTaoX fitting process, we first compute the maximum likelihood estimates (MLEs) of the model parameters, and use the them to initialize the Markov Chain Monte Carlo (MCMC) sampler for posterior sampling. The adopted MCMC priors are listed in Table 2 and are fixed throughout the analysis unless stated otherwise. The prior ranges are chosen according to the light curve properties. For example, the lower limit for $\tau_{\rm DRW}$ is set to $\sim$10 times smaller than the minimum time separation between any two observations, while the upper limit is $\sim$10 times larger than the light curve length. MCMC sampling is performed using the gradient-based No-U-Turn Sampler (NUTS; Hoffman & Gelman, 2011), with a single chain run for $10^{4}$ steps following $10^{3}$ warm-up steps. The posterior median of $\tau_{\rm DRW}$ is taken as the point estimate.
@@
We also apply JAVELIN to the first set of simulated light curves to measure $\tau_{\rm gr}$. The JAVELIN light curve model includes five parameters: $\tau_{\rm DRW}$, $\sigma_{\mathrm{DRW}}$, $s_{\rm gr}$, $w_{\rm gr}$, and $\tau_{\rm gr}$. Here, $s_{\rm gr}$ and $w_{\rm gr}$ describe the scale and width, respectively, of the top-hat transfer function for the ggrr lag. JAVELIN employs a two-step fitting procedure: it first generates a distribution of $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ by fitting the light curve of a chosen band via MCMC sampling, and then uses that distribution as a prior for the DRW parameters while simultaneously fitting both the DRW parameters and the interband lags from the light curves in all bands. In the first step, the light curve in the bluest band is typically used, as JAVELIN assumes that redder-band light curves are convolved versions of it; here we use the $g$-band light curve. For MCMC sampling, JAVELIN employs the affine-invariant ensemble sampler (Goodman & Weare, 2010) as implemented in emcee (Foreman-Mackey et al., 2013). In the first step we follow the default configuration, while in the second step we run the sampler with 100 walkers, 500 burn-in steps, and 3000 sampling steps. We also set the prior bound for $\tau_{\rm DRW}$ to $10^{-4}$, $10^{4}$] days and for $\tau_{\rm gr}$ to [15-15, $25$] days. The posterior median of $\tau_{\rm gr}$ is reported as the final point estimate.
@@
We clean the sample of best-fit EzTaoX and JAVELIN lags by removing those associated with non-convergent MCMC chains. Specifically, we discard fits in which the MCMC chain for $\tau_{\rm gr}$ yields fewer than 400 effective samples ($n_{\rm eff}$), following the recommendation of Vehtari et al. (2021) for selecting converged chains. Approximately 20% of the 3-year and 7% of the 10-year EzTaoX fits are removed by this criterion, while $\sim$20% of the JAVELIN fits are removed from both the 3-year and 10-year data. Examined as a function of the input $\tau_{\rm gr}$, the fraction of discarded JAVELIN fits increases with increasing true $\tau_{\rm gr}$; in contrast, the fraction of discarded EzTaoX fits remains relatively constant across the range of input $\tau_{\rm gr}$. We discuss this effect in more detail in Section 4.3.
@@ 3.1.3 Recovery of Stochastic Variability Timescale
To better demonstrate the advantages of multiband fitting, we refit the second set of simulated light curves after removing revisits occurring within 1 hr in any given band. This removal mimics a scenario in which intra-night observations must be stacked to detect faint variable sources against host-galaxy starlight—such as low-luminosity and low-mass AGNs, which are expected to exhibit shorter characteristic variability timescales than typical quasars (e.g., Burke et al., 2021). Figure 4 displays the recovered $\tau_{\rm DRW}$ as a function of the input value. When using only $u$- or $g$-band observations, the recovered $\tau_{\rm DRW}$ is systematically biased for input values below a few days. In contrast, combining $u$- and $g$-band data allows $\tau_{\rm DRW}$ to be recovered without bias down to hourly timescales. Furthermore, MCMC chains fail to converge (i.e., $n_{\rm eff}<400$) in more than 40% of the single-band fits, whereas this fraction decreases to only 7% for the multiband fits, underscoring the improved reliability of multiband fitting. The multiband capability of EzTaoX thus enables robust selection of low-luminosity and low-mass AGNs based on variability, particularly those with intrinsically faint emission and stochastic variability characterized by $\tau_{\rm DRW}$ values of just hours to days.
@@ 3.1.4 Recovery of Stochastic Variability Amplitude
Figure 6: Comparison of interband continuum lags for NGC 5548 measured with EzTaoX to those reported in the literature. The light curves of NGC 5548 are obtained by the AGN STORM project. In the legend, F16 refers to the lags reported by Fausnaugh et al. (2016), and E15 refers to the lags reported by Edelson et al. (2015). Interband continuum lags measured with EzTaoX are broadly consistent with those obtained using JAVELIN and the ICCF method, highlighting the effectiveness of EzTaoX despite its adoption of a simplified transfer function (i.e., a Dirac delta function).
Figure 7: Comparison of interband continuum lags for a sample of type 1 AGNs measured with EzTaoX to those reported in the literature (Guo et al., 2022). The light curves are obtained from ZTF. Interband continuum lags measured with EzTaoX are broadly consistent with those obtained using JAVELIN (left) and the ICCF method (right)
@@ 3.2 A Test Using Continuum Reverberation Mapping Light Curves of NGC 5548
The AGN Space Telescope and Optical Reverberation Mapping (STORM) project is an intensive multiwavelength monitoring campaign targeting NGC 5548, a well-studied Seyfert galaxy (De Rosa et al., 2015; Edelson et al., 2015; Fausnaugh et al., 2016). This campaign combines observations from the Hubble Space Telescope (HST), the Neil Gehrels Swift Observatory (Swift hereafter; Gehrels et al., 2004), and a network of ground-based telescopes, to provide simultaneous daily light curves of NGC 5548 across the X-ray, UV, and optical bands. For further details on the observational setup and data, we refer interested readers to De Rosa et al. (2015), Edelson et al. (2015), and Fausnaugh et al. (2016).
Applying EzTaoX to the UV/optical AGN STORM light curves of NGC 5548 reveals that the lags inferred by EzTaoX are broadly consistent with literature values derived using JAVELIN and ICCF. We follow the same strategy described in Section 3.1.2 to measure interband lags relative to the continuum light curve at 1367Å. In contrast to the approaches taken by Edelson et al. (2015), and Fausnaugh et al. (2016), we do not de-trend the light curves prior to lag estimation. We also increased the prior range of $S_{\rm band}$ to Uniform($^{-5}, 1010$) because AGN STORM light curve are provided in flux, which exhibits large amplitude variation across bands.
Figure 6 compares the lags measured with EzTaoX to those derived with JAVELIN and ICCF. F16 and E15 in the legend refer to the lags reported by Fausnaugh et al. (2016) and Edelson et al. (2015), respectively. The blue points represent the continuum lags measured at 1158Å, 1479Å, and 1746Å  using HST light curves. The green points correspond to the continuum lags from the Swift UVW2, UVM2, UVW1, U, and B bands. The yellow points represent continuum lags from Johnson/Cousins BVRI bands and Sloan Digital Sky Survey (SDSS; York et al., 2000) ugrizugriz bands based on ground-based observations. The purple squares in the right panel of Figure 6 shows the Swift lags reported by Edelson et al. (2015). We note that the Swift ICCF lags from Fausnaugh et al. (2016) are systematically larger than those reported by Edelson et al. (2015), likely due to the different de-trending approaches they adopted. Overall, interband lags measured using EzTaoX are broadly consistent with those obtained using JAVELIN and ICCF.
@@ 3.3 A Test Using Zwicky Transient Facility AGN Light Curves
We apply EzTaoX to the $g$- and $r$-band ZTF light curves of the core-sample AGNs from Guo et al. (2022), and find that the lags measured with EzTaoX are broadly consistent with those obtained using JAVELIN and ICCF. Figure 7 compares the interband lags measured with EzTaoX to those from JAVELIN (left panel) and ICCF (right panel). We retrieve the ZTF light curves for these sources from the IPAC Infrared Science Archive (ZTF Team, 2025) and truncate them to match the DR7 temporal baseline (i.e., MJD < 59400). When multiple light curves are available for the same object (each obtained from a different CCD), we select the one with the largest number of epochs. We follow the same strategy described in Section 3.1.2 to measure the lags between the $g$ and $r$ bands. We also fit additional jitter terms—one for each band—added to the diagonal elements of the covariance matrix to account for any residual photometric uncertainties not captured by the photometric pipeline (Burke et al., 2021).
@@ 4.1 Unbiased Point Estimator of $\tau_{\rm DRW}$
Unbiased recovery of $\tau_{\rm DRW}$ depends on the light-curve cadence, baseline (i.e., the total duration of observations), and adopted priors. As shown by Kozłowski (2017) and later studies (e.g., Sánchez-Sáez et al., 2018; Burke et al., 2021; Suberlak et al., 2021; Hu et al., 2024), $\tau_{\rm DRW}$ is systematically underestimated when the light-curve baseline is shorter than $\sim$10 times the true $\tau_{\rm DRW}$. Hu et al. (2024) conducted extensive simulations demonstrating that the assumed prior on $\tau_{\rm DRW}$ also impact its recovery. In addition, they reported that $\tau_{\rm DRW}$ is systematically overestimated when the true $\tau_{\rm DRW}$ is less than five times the mean cadence, since the temporal sampling then probes only the white-noise regime (i.e., a PL exponent of zero) of the DRW PSD.
@@ 4.2 Deviation from DRW
In this work, we assume that AGN stochastic variability follows a DRW process; however, the modeling framework introduced here is sufficiently flexible to accommodate alternative GP kernels for characterizing the underlying variability. As an illustration, Figure 10 presents $g$- and $r$-band light curves simulated using a CARMA(2,1) process, sampled according to the LSST WFD cadence (left), along with the corresponding PSDs recovered using EzTaoX (right). The CARMA(2,1) process—also known as the noise-driven damped harmonic oscillator (DHO) model—has been shown to provide superior fits to observed AGN UV/optical light curves compared to the DRW model (e.g., Kasliwal et al., 2017; Yu et al., 2025). For reference, the DRW PSD is shown in blue in the right panel. Similar to Figure 1, the orange histogram in the lower right corner displays the posterior distribution of the ggrr lag. The specific CARMA(2,1) process simulated is governed by the following kernel:

### v2

@@ 3.1.1 Light Curve Simulation
We adopt the DRW model to simulate AGN light curves in individual LSST bands. We start by simulating DRW light curves in each LSST band with a 30-second cadence; this 30-second cadence is chosen to match the 30-second exposure time of each LSST single visit (ivezic2019). The input DRW parameters across different bands are correlated according to the wavelength-dependent power-law relations presented in macleod2010. More specifically,
@@
where $\lambda$ is the effective wavelength of the LSST filters (ivezic2019). We then simulate continuum lags by offsetting light curves in redder band in time with respect to light curves in bluer bands. Finally, we down-sample the simulated 30-second cadence light curve to the observing cadence of the WFD survey.
@@
where $\sigma_{\mathrm{sys}}$ is the systematic photometric error and $\sigma_{\mathrm{rand}}$ is the random photometric error. LSST aims to achieve a $\sigma_{\mathrm{sys}}<0.005$ mag (ivezic2019); we set $\sigma_{\mathrm{sys}}$ to $0.005$ mag in our simulation. The random photometric error ($\sigma_{\mathrm{rand}}$) is expected to scale as
@@
The observing cadence and the expected $m_{5}$ at any location on the sky are predicted by LSST operation simulations (OpSims) using MAF (marshall2017). In this work, we adopt the operation simulation version baseline_v4.0_10yrs. All our simulated AGNs also assume a fiducial $g$-band mean magnitude of 21 ($\sim 1\%$ photometric accuracy) and mean uugg and ggrr colors of 0.1 mag (richards2001; temple2021). An example light curve from the first set is shown in Figure 1.
@@
The first set aims to evaluate EzTaoX’s ability to recover interband lags (see Section 3.1.2), and the light curves consist of simulated LSST observations in $g$ and $r$ bands. The ggrr lag for each multiband light curve is uniformly sampled between 0.01 day and 10 days, while the input $g$-band $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ are fixed at fiducial values of 100 days and 0.112 mag, respectively. These choices of $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ are representative of those measured from typical quasar (i.e., high luminosity AGN) UV/optical light curves (e.g., vandenberk2004; macleod2010).
@@ 3.1.2 Recovery of Interband Lag
We run EzTaoX on the first set of simulated light curves to extract the lag between the $g$- and $r$-band ($\tau_{\rm gr}$), after subtracting the mean from each single-band light curve. The EzTaoX light curve model includes seven parameters: $\tau_{\rm DRW}$, $\sigma_{\mathrm{DRW}}$, $S_{\rm g}$, $S_{\rm r}$, $\mu_{\rm g}$, $\mu_{\rm r}$, and $\tau_{\rm gr}$. Here, $S_{\rm g}$ and $S_{\rm r}$ are the GP kernel scaling factors for $g$- and $r$-band observations (see Equation 5), while $\mu_{\rm g}$ and $\mu_{\rm r}$ are the mean magnitudes of the $g$- and $r$-band light curves, respectively. As part of the EzTaoX fitting process, we first compute the maximum likelihood estimates (MLEs) of the model parameters, and use the them to initialize the Markov Chain Monte Carlo (MCMC) sampler for posterior sampling. The adopted MCMC priors are listed in Table 2 and are fixed throughout the analysis unless stated otherwise. The prior ranges are chosen according to the light curve properties. For example, the lower limit for $\tau_{\rm DRW}$ is set to $\sim$10 times smaller than the minimum time separation between any two observations, while the upper limit is $\sim$10 times larger than the light curve length. MCMC sampling is performed using the gradient-based No-U-Turn Sampler (NUTS; Hoffman2011), with a single chain run for $10^{4}$ steps following $10^{3}$ warm-up steps. The posterior median of $\tau_{\rm DRW}$ is taken as the point estimate.
@@
We also apply JAVELIN to the first set of simulated light curves to measure $\tau_{\rm gr}$. The JAVELIN light curve model includes five parameters: $\tau_{\rm DRW}$, $\sigma_{\mathrm{DRW}}$, $s_{\rm gr}$, $w_{\rm gr}$, and $\tau_{\rm gr}$. Here, $s_{\rm gr}$ and $w_{\rm gr}$ describe the scale and width, respectively, of the top-hat transfer function for the ggrr lag. JAVELIN employs a two-step fitting procedure: it first generates a distribution of $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ by fitting the light curve of a chosen band via MCMC sampling, and then uses that distribution as a prior for the DRW parameters while simultaneously fitting both the DRW parameters and the interband lags from the light curves in all bands. In the first step, the light curve in the bluest band is typically used, as JAVELIN assumes that redder-band light curves are convolved versions of it; here we use the $g$-band light curve. For MCMC sampling, JAVELIN employs the affine-invariant ensemble sampler (goodman2010) as implemented in emcee (foreman-mackey2013). In the first step we follow the default configuration, while in the second step we run the sampler with 100 walkers, 500 burn-in steps, and 3000 sampling steps. We also set the prior bound for $\tau_{\rm DRW}$ to $10^{-4}$, $10^{4}$] days and for $\tau_{\rm gr}$ to [15-15, $25$] days. The posterior median of $\tau_{\rm gr}$ is reported as the final point estimate.
@@
We clean the sample of best-fit EzTaoX and JAVELIN lags by removing those associated with non-convergent MCMC chains. Specifically, we discard fits in which the MCMC chain for $\tau_{\rm gr}$ yields fewer than 400 effective samples ($n_{\rm eff}$), following the recommendation of Vehtari2021 for selecting converged chains. Approximately 20% of the 3-year and 7% of the 10-year EzTaoX fits are removed by this criterion, while $\sim$20% of the JAVELIN fits are removed from both the 3-year and 10-year data. Examined as a function of the input $\tau_{\rm gr}$, the fraction of discarded JAVELIN fits increases with increasing true $\tau_{\rm gr}$; in contrast, the fraction of discarded EzTaoX fits remains relatively constant across the range of input $\tau_{\rm gr}$. We discuss this effect in more detail in Section 4.3.
@@ 3.1.3 Recovery of Stochastic Variability Timescale
To better demonstrate the advantages of multiband fitting, we refit the second set of simulated light curves after removing revisits occurring within 1 hr in any given band. This removal mimics a scenario in which intra-night observations must be stacked to detect faint variable sources against host-galaxy starlight—such as low-luminosity and low-mass AGNs, which are expected to exhibit shorter characteristic variability timescales than typical quasars (e.g., burke2021). Figure 4 displays the recovered $\tau_{\rm DRW}$ as a function of the input value. When using only $u$- or $g$-band observations, the recovered $\tau_{\rm DRW}$ is systematically biased for input values below a few days. In contrast, combining $u$- and $g$-band data allows $\tau_{\rm DRW}$ to be recovered without bias down to hourly timescales. Furthermore, MCMC chains fail to converge (i.e., $n_{\rm eff}<400$) in more than 40% of the single-band fits, whereas this fraction decreases to only 7% for the multiband fits, underscoring the improved reliability of multiband fitting. The multiband capability of EzTaoX thus enables robust selection of low-luminosity and low-mass AGNs based on variability, particularly those with intrinsically faint emission and stochastic variability characterized by $\tau_{\rm DRW}$ values of just hours to days.
@@ 3.1.4 Recovery of Stochastic Variability Amplitude
Figure 6: Comparison of interband continuum lags for NGC 5548 measured with EzTaoX to those reported in the literature. The light curves of NGC 5548 are obtained by the AGN STORM project. In the legend, F16 refers to the lags reported by fausnaugh2016, and E15 refers to the lags reported by Edelson2015. Interband continuum lags measured with EzTaoX are broadly consistent with those obtained using JAVELIN and the ICCF method, highlighting the effectiveness of EzTaoX despite its adoption of a simplified transfer function (i.e., a Dirac delta function).
Figure 7: Comparison of interband continuum lags for a sample of type 1 AGNs measured with EzTaoX to those reported in the literature (guo2022). The light curves are obtained from ZTF. Interband continuum lags measured with EzTaoX are broadly consistent with those obtained using JAVELIN (left) and the ICCF method (right)
@@ 3.2 A Test Using Continuum Reverberation Mapping Light Curves of NGC 5548
The AGN Space Telescope and Optical Reverberation Mapping (STORM) project is an intensive multiwavelength monitoring campaign targeting NGC 5548, a well-studied Seyfert galaxy (DeRosa2015; Edelson2015; fausnaugh2016). This campaign combines observations from the Hubble Space Telescope (HST), the Neil Gehrels Swift Observatory (Swift hereafter; Gehrels2004), and a network of ground-based telescopes, to provide simultaneous daily light curves of NGC 5548 across the X-ray, UV, and optical bands. For further details on the observational setup and data, we refer interested readers to DeRosa2015, Edelson2015, and fausnaugh2016.
Applying EzTaoX to the UV/optical AGN STORM light curves of NGC 5548 reveals that the lags inferred by EzTaoX are broadly consistent with literature values derived using JAVELIN and ICCF. We follow the same strategy described in Section 3.1.2 to measure interband lags relative to the continuum light curve at 1367Å. In contrast to the approaches taken by Edelson2015, and fausnaugh2016, we do not de-trend the light curves prior to lag estimation. We also increased the prior range of $S_{\rm band}$ to Uniform($^{-5}, 1010$) because AGN STORM light curve are provided in flux, which exhibits large amplitude variation across bands.
Figure 6 compares the lags measured with EzTaoX to those derived with JAVELIN and ICCF. F16 and E15 in the legend refer to the lags reported by fausnaugh2016 and Edelson2015, respectively. The blue points represent the continuum lags measured at 1158Å, 1479Å, and 1746Å  using HST light curves. The green points correspond to the continuum lags from the Swift UVW2, UVM2, UVW1, U, and B bands. The yellow points represent continuum lags from Johnson/Cousins BVRI bands and Sloan Digital Sky Survey (SDSS; york2000) ugrizugriz bands based on ground-based observations. The purple squares in the right panel of Figure 6 shows the Swift lags reported by Edelson2015. We note that the Swift ICCF lags from fausnaugh2016 are systematically larger than those reported by Edelson2015, likely due to the different de-trending approaches they adopted. Overall, interband lags measured using EzTaoX are broadly consistent with those obtained using JAVELIN and ICCF.
@@ 3.3 A Test Using Zwicky Transient Facility AGN Light Curves
We apply EzTaoX to the $g$- and $r$-band ZTF light curves of the core-sample AGNs from guo2022, and find that the lags measured with EzTaoX are broadly consistent with those obtained using JAVELIN and ICCF. Figure 7 compares the interband lags measured with EzTaoX to those from JAVELIN (left panel) and ICCF (right panel). We retrieve the ZTF light curves for these sources from the IPAC Infrared Science Archive (ztf_lightcurve) and truncate them to match the DR7 temporal baseline (i.e., MJD < 59400). When multiple light curves are available for the same object (each obtained from a different CCD), we select the one with the largest number of epochs. We follow the same strategy described in Section 3.1.2 to measure the lags between the $g$ and $r$ bands. We also fit additional jitter terms—one for each band—added to the diagonal elements of the covariance matrix to account for any residual photometric uncertainties not captured by the photometric pipeline (burke2021).
@@ 4.1 Unbiased Point Estimator of $\tau_{\rm DRW}$
Unbiased recovery of $\tau_{\rm DRW}$ depends on the light-curve cadence, baseline (i.e., the total duration of observations), and adopted priors. As shown by kozlowski2017a and later studies (e.g., sanchez-saez2018; burke2021; suberlak2021; Hu2024), $\tau_{\rm DRW}$ is systematically underestimated when the light-curve baseline is shorter than $\sim$10 times the true $\tau_{\rm DRW}$. Hu2024 conducted extensive simulations demonstrating that the assumed prior on $\tau_{\rm DRW}$ also impact its recovery. In addition, they reported that $\tau_{\rm DRW}$ is systematically overestimated when the true $\tau_{\rm DRW}$ is less than five times the mean cadence, since the temporal sampling then probes only the white-noise regime (i.e., a PL exponent of zero) of the DRW PSD.
@@ 4.2 Deviation from DRW
In this work, we assume that AGN stochastic variability follows a DRW process; however, the modeling framework introduced here is sufficiently flexible to accommodate alternative GP kernels for characterizing the underlying variability. As an illustration, Figure 10 presents $g$- and $r$-band light curves simulated using a CARMA(2,1) process, sampled according to the LSST WFD cadence (left), along with the corresponding PSDs recovered using EzTaoX (right). The CARMA(2,1) process—also known as the noise-driven damped harmonic oscillator (DHO) model—has been shown to provide superior fits to observed AGN UV/optical light curves compared to the DRW model (e.g., kasliwal2017; Yu2025). For reference, the DRW PSD is shown in blue in the right panel. Similar to Figure 1, the orange histogram in the lower right corner displays the posterior distribution of the ggrr lag. The specific CARMA(2,1) process simulated is governed by the following kernel: